ent Exercise 8-5 has accounts receivable of $97.900 at March 31, 2017, Credit terms are 2/10, n/30. At March 31, 2017, there is a $1,951 credit balance in Doubtful Accounts prior to adjustment. The company uses the percentage-of-receivables shown below. basis for estimating uncollectible accounts.T ,951 credit balance in Allowance for estimates of bad debts are as Balance, March 31 Estimated Percentage Current $69,000 $80,200 1-30 days past due 12,000 7,260 31-90 days past due 9,900 2,500 Over 90 days past due 7.000 1,060 97,900 $91,020 7% 28 53 Determine the total estimated uncollectibles. The total estimated uncollectibles Prepare the adjusting entry at March 31, 2017, to record bad debt ex Credit account titles are automati ally indented when amount斥onternd. Do not in ant pense. Arcount Tirlos and Fnlanation nobit Credit
The total estimated uncollectibles SNOW 11ST OF ACCOUNTS LINK TO TEXT VIDEO: APPLIED SKILLS VIDEO: SIMILAR EXERCISE Prepare the adjusting entry at March 31, 2017, to record bad debt expense. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit LIST OF ACCOUNTS LINK T0 TEX VIDEO: APPLİID SKILLS VIDEO SIMILAR EXERCISE Question Attempts: 1 of 3 used SAVE FOR LATER SUBHIT ANSWER
Answer #1

Total estimated uncollectibles=69000*2%+12000*5%+9900*28%+7000*53%

=1380+600+2772+3710=8462

Assumption only the balance in the year 2017 is taken for calculating the % of uncollectibles

Already the company has a bad debt allowance of $1951cr so the difference 8462-1951=6511 is the actual bad debt expense

Journal entry

Bad debts expense a/c Dr. 6511

To Allowance for bad debts Cr 6511